OpenStudy (anonymous):

the worlds oldest bride was 19 yr older than her groom. together, their ages totaled 185yr. how old were the bride and the groom.

OpenStudy (anonymous):

i could only think of 100 and 85

OpenStudy (anonymous):

Chose variables for the age of the bride and for the age of the groom then write equations that relate them

OpenStudy (anonymous):

i get it thx

OpenStudy (anonymous):

wait i think i did it wrong

OpenStudy (anonymous):

my answer is wrong i can tell cant be 97 and

Directrix (directrix):

Let G represent the groom's age. Then the bride's age, B, is 19 years more. Let B = 19 + G

Directrix (directrix):

G + B = 185 G + (19 + G) = 185 2G + 19 + 185 Solve for G Add 19 to the G value to get the B value. @cr7